OpenStudy (anonymous):

Find the point P where the line x = 1 + t, y = 2t, z = -3t intersects the plane x + y - z = 1.

OpenStudy (helder_edwin):

just plug in the "values" for x, y and z into the plane's equation. then solve for t

OpenStudy (anonymous):

whar values?

OpenStudy (anonymous):

oh nevermind i get it, thanks

OpenStudy (helder_edwin):

x=1+t y=2t z=-3t then \[ \large 1=x+y-z=(1+t)+(2t)-(-3t) \]